OpenStudy (anonymous):

The length of a rectangle is 5m longer than its width. If the perimeter of the rectangle is 42m, find its area.

OpenStudy (nurali):

First, you know that the length of the rectangle is 5m longer than the width, which can be written as: L = W + 5 (with W=width and L= length) Secondly,you know that a rectangle has 4 sides with each of the opposite sides having the same length, and the total perimeter is 42m. So: W+W+L+L = 42 or 2W + 2L = 42 Now you can plug the first equation into the second equation, replacing the L value: 2W + 2L = 42 2W + 2(W+5) = 42 2W + 2W + 10 = 42 4W + 10 = 42 4W = 32 W = 8 Now plug this into the first equation to find the length: L = W + 5 L = 8+5 L = 13 So the length of the rectangle is 13m and the width is 8m. Finally, use the area of a rectangle equation (A=L*W) to finish the problem: A = L*W A = 13*8 A = 104 meters squared So the area of the rectangle is 104 meters squared.